40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 64 of 327  Not logged in ELOG logo
ID Date Author Type Category Subject
  13290   Mon Sep 4 18:18:29 2017 ranaUpdateLSCdewhite switching: FOTON settings

not immediately necessary, since you have already got it sort of working, but one of these days we should optimize this for real. In the past, we used to do this by putting a o'scope on the coil Vmon during the switching to catch the transient w/ triggering. We download the data/picture via ethernet. Run for loop on tolerance to see what's what.

  1. Went into the Foton filter banks for all the coil output filters, and modified the "Output" settings to be on "Input crossing", with a "Tolerance" of 10 and a "Timeout" of 3 seconds. These settings are to facilitate smooth transition between the two signal paths (without and with coil-dewhitening). The parameters chosen were arbitrary and not optimized in any systematic manner.

 

  13289   Mon Sep 4 16:30:06 2017 gautamUpdateLSCOplev loop tweaking

Now that the DRMI locking seems to be repeatable again, I want to see if I can improve the measured MICH noise. Recall that the two dominant sources of noise were

  1. BS Oplev loop A2L - this was the main noise between 30-60Hz.
  2. DAC noise - this dominated between ~60-300Hz, since we were operating with the de-whitening filters off.

In preparation for some locking attempts today evening, I did the following:

  1. Added steeper elliptic roll-off filters for the ITMX and ITMY Oplevs. This is necessary to allow the de-whitening filters to be turned on without railing the DAC.
  2. Modified the BS Oplev loop to also have steeper high-frequency (>30Hz) roll off. The roll-off between 15-30Hz is slightly less steep as a result of this change.
  3. Measured all Oplev loop TFs - UGFs are between 4 Hz and 5 Hz, phase margin is ~30degrees. I did not do any systematic optimization of this for today.
  4. Went into the Foton filter banks for all the coil output filters, and modified the "Output" settings to be on "Input crossing", with a "Tolerance" of 10 and a "Timeout" of 3 seconds. These settings are to facilitate smooth transition between the two signal paths (without and with coil-dewhitening). The parameters chosen were arbitrary and not optimized in any systematic manner.
  5. After making the above changes, I tried engaging the de-whitening filters on ITMX, ITMY and BS with the arms locked. In the past, I was unable to do this because of a number of issues - Oplev loop shapes and Foton settings among them. But today, the switching was smooth, the single arm locks weren't disturbed when I engaged the coil de-whitening.

Hopefully, I can successfully engage a similar transition tonight with the DRMI locked. The main difference compared to this daytime test is going to be that the MICH control signal is also going to be routed to the BS.

Tasks for tonight, if all goes well:

  1. Lock DRMI.
  2. Use UGF servos to set the overall loop gains for DRMI DoFs.
  3. Reduce PRCL->MICH and SRCL->MICH coupling.
  4. Measure loop shapes of all DRMI DoFs.
  5. Make sensing matrix measurement.
  6. Engage coil-dewhitening, download data, make NB.

Unrelated to this work: the PMC was locked near the upper rail of the PZT, so I re-locked it closer to the middle of the range.

Quote:

Surprisingly, there was no evidence of REFL55 behaving weirdly tonight, and I was able to easily lock the DRMI on 1f error signals using the recipe I've been using in the last few months.

  13288   Fri Sep 1 19:15:40 2017 gautamUpdateALSFiber ALS noise measurement

Summary:

I did some work today to see if I could use the IR beat for ALS control. Initial tests were encouraging.

I will now embark on the noise budgeting.

Details:

  • For this test, I used the X arm
  • I hooked up the X-arm + PSL IR beat to the X-arm DFD channel, and used the Y-arm DFD channels to simultaneously monitor the X-arm green beat.
  • I then transitioned to ALS control and used POX as an out-of-loop sensor for the ALS noise.
  • Attachment #1 shows a comparison of the measurements. In red is the IR beat, while the green traces are from the test EricQ and I did a couple of nights ago using the green beat.
  • I also wanted to do some arm cavity scans with the arm under ALS control with the IR beat - but was unsucessful. The motivation was to fix the ALS model counts->Hz calibration factors.
  • I did however manage to do a 10 FSR scan using the green beatnote - however, towards the end of this scan, the green beat frequency (read off the control room analyzer) was ~140MHz, which I believe is outside (or at least on the edge) of the bandwidth of the Green BBPDs. The fiber coupled IR beat photodiodes have a much larger (1GHz) spec'd bandwidth.

I am leaving the green beat electronics on the PSL table in the switched state for further testing...

 

Attachment 1: IR_ALS_noise.pdf
IR_ALS_noise.pdf
  13287   Fri Sep 1 16:55:27 2017 gautamUpdateComputersTestpoints now accessible again

Thanks to Jonathan Hanks, it appears we can now access test-points again using dataviewer.

I haven't done an exhaustive check just yet, but I have loaded a few testpoints in dataviewer, and ran a script that use testpoint channels (specifically the ALS phase tracker UGF setting script), all seems good.

So if I remember correctly, the major CDS fix now required is to solve the model unloading issue.

Thanks to Jamie/Jonathan Hanks/KT for getting us back to this point! Here are the details:

After reading logs and code, it was a simple daqdrc config change.

The daqdrc should read something like this:

...
set master_config=".../master";
configure channels begin end;
tpconfig ".../testpoint.par";
...


What had happened was tpconfig was put before the configure channels
begin end.  So when daqd_rcv went to configure its test points it did
not have the channel list configured and could not match test points to
the right model & machine.  Dave and I suspect that this is so that it
can do an request directly to the correct front end instead of a general
broadcast to all awgtpman instances.

Simply reordering the config fixes it.

I tested by opening a test point in dataviewer and verifiying that
testpoints had opened/closed by using diag -l.  Xmgr/grace didn't seem
to be able to keep up with the test point data over a remote connection.

You can find this in the logs by looking for entries like the following
while the daqd is starting up.  When we looked we saw that there was an
entry for every model.

Unable to find GDS node 35 system c1daf in INI fiels
  13286   Fri Sep 1 16:27:39 2017 gautamUpdateSUSMC1 glitching

I re-enabled the MC SUS damping and IMC locking for some IFO work just now.

Quote:

MC1, MC2 and MC3 damping turned off to see glitching action at 9:57am

 

  13285   Fri Sep 1 15:46:12 2017 KiraUpdatePEMtemp sensor update

I took off the AD590 and attached it to two long wires leading out from the board. This will allow us to attach the sensor to a metal block and not have to stick the whole board to it. I have also completed three identical copies of this and it's pretty much ready to be tested. According to Craig and Andrew's elog here, the sensor is very noisy and they added in a low pass filter to fix that, so that's something to consider for the final version of the circuit. I'll test what I have so far and see how that goes. We still need to figure out how to get readings from the sensors.

To attach the sensor to the metal block, I'll use some thermal paste and fasteners. I'll also put a thermometer on the block to record the actual temperature. I'll then wrap it in some insulation we have in the lab and have only some wires leading out of it to make measurements. I'll leave this setup overnight and record the outputs for about a full day. The fluctuations between the sensors will then indicate the noise of each individual sensor.

Attachment 1: IMG_20170901_144729.jpg
IMG_20170901_144729.jpg
  13284   Fri Sep 1 08:25:08 2017 SteveUpdateSUSMC1 glitching

MC1, MC2 and MC3 damping turned off to see glitching action at 9:57am

Quote:

There was a pretty large glitch in MC1 about an hour ago. The misalignment was so large that the autolocker wasn't able to lock the IMC. I manually re-aligned MC1 using the bias sliders, and now IMC locks fine. Attached is a 90 second plot of 2K data from the OSEMs showing the glitch. Judging from the wall StripTool, the IMC was well behaved for ~4 hours before this glitch - there is no evidence of any sort of misalignment building up, judging from the WFS control signals.

 

Attachment 1: MC1glitching.png
MC1glitching.png
Attachment 2: MC1kicks.png
MC1kicks.png
  13283   Thu Aug 31 21:40:24 2017 gautamUpdateGeneralMC1 kicked again

There was a pretty large glitch in MC1 about an hour ago. The misalignment was so large that the autolocker wasn't able to lock the IMC. I manually re-aligned MC1 using the bias sliders, and now IMC locks fine. Attached is a 90 second plot of 2K data from the OSEMs showing the glitch. Judging from the wall StripTool, the IMC was well behaved for ~4 hours before this glitch - there is no evidence of any sort of misalignment building up, judging from the WFS control signals.

Attachment 1: MC1_glitch.png
MC1_glitch.png
  13282   Thu Aug 31 18:36:23 2017 gautamUpdateCDSrevisiting Acromag

Current status:

  • There is a single Acromag ADC unit installed in 1X4
  • It is presently hooked up to the PSL NPRO diagnostic connector channels
  • I had (re)-started the acquisiton of these channels on August 16 - but for reasons unknown, the tmux session that was supposed to be running the EPICS server on megatron seems to have died on August 22 (judging by the trend plot of these channels, see Attachment #1)
  • I had not set up an upstart job that restarts the server automatically in such an event. I manually restarted it for now, following the same procedure as linked in my previous elog.
  • While I was at it, I also took the opportunity to edit the Acromag channel names to something more appropriate - all channels previously prefixed with C1:ACRO- have now been prefixed with C1:PSL-

Plan of action:

  1. Hardware - we have, in the lab, in addition to the installed ADC unit
    • 3x 8 channel differential input ADC units
    • 2x 8 channel differential output DAC units
    • 1x 16 channel BIO unit
    • 2U chassis + connectors + breakout boards + other misc hardware that I think Johannes and Lydia procured with the original plan to replace the EX slow controls.
    • Some relevant elogs: Panel designs, breakout design, sketch for proposed layout, preliminary channel list.
      So on the hardware side, it would seem that we have everything we need to go ahead with replacing the EX slow controls with an Acromag system, although Johannes probably knows more about our state of readiness from a hardware PoV.
  2. Software
    • We probably want to get a dedicated machine that will handle the EPICS channel serving for the Acromag system
    • Have to figure out the networking arrangement for such a machine
    • Have to figure out how to set up the EPICS server protocol in such a way that if it drops for whatever reason, it is automatically restarted

 

Attachment 1: Acromag_EPICS.png
Acromag_EPICS.png
  13281   Thu Aug 31 03:31:15 2017 gautamUpdateLSCDRMI re-locked!

After our Demod/Whitening electronics investigations suggested nothing obviously wrong, I decided to give DRMI locking another go tonight.

Surprisingly, there was no evidence of REFL55 behaving weirdly tonight, and I was able to easily lock the DRMI on 1f error signals using the recipe I've been using in the last few months.

Not sure what to make of all this frown.

I got in a ~15 minute lock, but I wasn't prepared to do any sort of characterization/ sensing / attempt to turn on coil-dewhitening, and I'm too tired to try again tonight. I was however able to whiten the error signals, as I have been able to do in the past. There is a ~45Hz bump in MICH that I haven't seen in the past.

I'll try and do some characterization tomorrow eve, but it's encouraging to at least get back to the pre-FB-failure state of locking.

Attachment 1: DRMI_1f.png
DRMI_1f.png
Attachment 2: DRMI_relocked.pdf
DRMI_relocked.pdf
  13280   Thu Aug 31 00:52:52 2017 gautamUpdateLSCREFL55 whitening board debugging

[rana,gautam]

We did an ingenious checkup of the whitening board tonight.

  • The board is D990694
  • We made use of a tip-tilt DAC channel for this test (specifically TT1 UL, which is channel 1 on the AI board). We disconnected the cable going from the AI board to the TT coil driver board.
    • as opposed to using a function generator to drive the whitening filter, this approach allows us to not have to worry the changing offsets as we switch the whitening gain.
    • By using the CDS system to generate the signal and also demodulate it, we also don't have to worry about the drive and demod frequencies falling out of sync with each other.
  • The test was done by injecting a low frequency (75.13 Hz, amplitude=0.1) excitation to this DAC channel, and using the LSC sensing matrix infrastructure to demodulate REFL55 I and Q at this frequency. Demod phases in these servos were adjusted such that the Q phase demodulated signal was minimized.
  • An excitation was injected using awggui into TT1 UL exc channel.
  • We then stepped the whitening gains for REFL55_I and REFL55_Q in 3dB steps, waiting 5 seconds for each step. Syntax is z step -s 5 C1:LSC-REFL55_I_WhiteGain +1.0,15 C1:LSC-REFL55_Q_WhiteGain +1.0,15
  • Attachment #1 suggests that the whitening filter board is working as expected (each step is indeed 3dB and all steps are equal to the eye).
  • Data + script used to generate this plot is in Attachment #2.

I've restored all connections at that we messed with at the LSC rack to their original positions.

The TT alignment seems to be drifting around more than usual after we disconnected one of the channels - when I came in today afternoon, the spot on the AS camera had drifted by ~1 spot diameter so I had to manually re-align TT1. 

Quote:
 

Based on my tests, everything on the Demod board seems to work as expected. I need to think more about what else could be happening here - specifically do a more direct test on the whitening board.

Attachment 1: REFL55_whtCheck.pdf
REFL55_whtCheck.pdf
Attachment 2: REFL55_whtChk.tar.gz
  13279   Thu Aug 31 00:46:57 2017 ranaSummaryCDSallegra -> Scientific Linux 7.3

I made a 'LiveCD' on a 16 GB USB stick using this command after the GUIs didn't work and looking at some blog posts:

sudo dd if=SL-7.3-x86_64-2017-01-20-LiveCD.iso of=/dev/sdf

Quote:

Debian doesn't like EPICS. Or our XY plots of beam spots...Sad!

Quote:
Quote:

No, not confused on that point. We just will not be testing OS versions at the 40m or running multiple OS's on our workstations. As I've said before, we will only move to so-called 'reference' systems once they've been in use for a long time.

Ubuntu16 is not to my knowledge used for any CDS system anywhere.  I'm not sure how you expect to have better support for that.  There are no pre-compiled packages of any kind available for Ubuntu16.  Good luck, you big smelly doofuses. Nyah, nyah, nyah.

K Thorne recommends that we use SL7.3 with the 'xfce' window manager instead of the Debian family of products, so we'll try it out on allegra and rossa to see how it works for us. Hopefully the LLO CDS team will be the tip of the spear on solving the usual software problems we have when we "~up" grade.

  13278   Thu Aug 31 00:19:35 2017 rana[^r]UpdatePSLIMC/FSS FAST gain

nominal changed from 22 to 23 dB to minimize PC drive RMS

previous loop gain measurement is sort of bogus (made on SR785); need some 4395 loop measurements and checking of crossover and error point spectrum

  13277   Wed Aug 30 22:15:47 2017 ranaOmnistructureComputersUSB flash drives moved

I have moved the USB flash drives from the electronics bench back into the middle drawer of the cabinet next to the AC which is west of the fridge. Drawer re-enlabeled.

  13276   Wed Aug 30 19:49:33 2017 gautamUpdateLSCREFL55 demod board debugging

Summary:

Today I tried debugging the mysterious increase in REFL55 signal levels in the DRMI configuration. I focused on the demod board, because last week, I had tried routing these signals through different channels on the whitening board, and saw the same effect. 

Based on my tests, everything on the Demod board seems to work as expected. I need to think more about what else could be happening here - specifically do a more direct test on the whitening board.

Details:

  • The demod board is a modified D990511 (marked up schematic + high-res photo to follow).
  • Initially, I tried probing the LO signal levels at various points with the board in the eurocrate itself, with the help of an extender card.
  • But this wasn't very convenient, so I pulled the board out to the office area for more testing.
  • The 55MHz LO signal going into the board is ~0dBm (measured with Agilent network analyzer)
  • I used the active probe to check the LO levels at various points along the signal chain, which mostly consists of attenuators, ERA-5SM amplifiers, and some splitters/phase rotators.
  • Everything seemed consistent with the expected levels based on "typical" numbers for gains and insertion losses cited in the datasheets for these devices.
  • I couldn't directly measure the level at the LO input to the mixer, but measuring the input to the ERA-5SM immediately before the mixer, barring problems with this amplifier, the LO input of the mixer is being driven at >17dBm which is what it wants.
  • Next, I decided to check the gain, gain imbalance and orthogonality of the demodulation.
  • For this purpose, I restored the board to the Eurocrate, reconnected the LO input to the board, and used a second Marconi at a slightly offset frequency to drive the PD input at ~0dBm.
  • Attachment #1 - The measured outputs look pretty balanced and orthogonal. The gain is consistent with an earlier measurement I made some months ago, when things were "normal". More bullets added after Rana's questions:
    • 300 MHz bandwidth oscilloscope used to acquire the data
    • I and Q outputs were from the daughter board
    • Data was acquired via ethernet data download utility
    • 20 MHz low-pass filter turned on on the Oscilloscope while downloading the data
Quote:

I did a quick check by switching the output of the REFL55 demod board to the inputs normally used by AS55 signals on the whitening board. Setting the whitening gain to +18dB for these channels had the same effect - ADC overflow galore. So looks like the whitening board isn't to blame. I will have to check the demod board out.

 


All connections have been restored untill further debugging later in the evening.

Attachment 1: REFL55_demod_check.pdf
REFL55_demod_check.pdf
  13275   Wed Aug 30 15:00:06 2017 gautamUpdateGeneralEdgeswitch fiber swap

A couple of minutes ago, Larry W swapped the fibers to our 40m Edgeswitch (BROCADE FWS 648G) to a faster connection. This is the switch to which our gateway machine, NODUS, is connected. The actual swap itself happened at the core router in Bridge, and took only a few seconds. After the switch, I double checked that I was able to ssh into nodus from my laptop, and Larry informed me that everything is working as expected on his end.

Larry also tells us that the other edgeswitch at the 40m (Foundry Networks), to which most of our GC network machines are connected, is a 100MBPS switch, and so we should re-route the connections from this switch to the BROCADE switch at our convenience to take advantage of the faster connection.

  13274   Wed Aug 30 11:04:08 2017 GabrieleSummaryLSCFirst look at neural network reconstruction of PRMI motion

Introduction

I trained a deep neural network (DNN) to reconstruct MICH and PRCL degrees of freedom in the PRMI configuration. For details on the DNN architecture please refer to G1701455 or G1701589. Or if you really want all the details you can look at the code. I used the following signals as input to the DNN: POPDC, POP22_Q, ASDC, REFL11_I/Q, REFL55_I/Q, AS55_I/Q.

Gautam took some PRMI data in free swinging and driven configuration:

  • 1187819331 + 10mins: Free swinging PRMI (after first locking PRMI on carrier and dither aligning).
  • 1187820070 + 5mins: PRM driven at low freq.
  • 1187820446 + 5mins: BS driven at low freq.

In contrast to the Fabry-Perot cavity case, we don't have a direct measurement of the real PRCL/MICH degrees of freedom, so it's more difficult to assess if the DNN is working well.

Results

All MICH and PRCL values are wrapped into the unique region [-lambda/4, lambda/4]^2. It's even a bit more complicated than simpling wrapping. Indeed, MICH is periodic over [-lambda/2, lambda/2]. However, the Michelson interferometer reflectivity (as seen from PRC) in the first half of the segment is  the same as in the second half, except for a change in sign. This change of sign in Michelson reflectivity can be compensated by moving PRCL by lambda/4, thus generating a pi phase shift in the PRC round trip propagation that compensate for the MICH sign change. Therefore, the unit cell of unique values for all signals can be taken as [-lambda/4, lambda/4] x [-lambda/4, lambda/4] for MICH x PRCL. But when we hit the border of the MICH region, PRCL is also affected by addtion of lambda/4. Graphically, the square regions A B C below are all equivalent, as well as more that are not highlighted:

This makes it a bit hard to un-wrap the resonstructed signal, especially when you add in the factor that in the reconstruction the wrapping is "soft".

The plot below shows an example of the time domain reconstruction of MICH/PRCL during the free swinging period.

It's hard to tell if the positions look reasonable, with all the wrapping going on.

Two-dimensional maps of signals

Here's an attempt at validating the DNN reconstruction. Using the reconstructed MICH/PRCL signal, I can create a 2d map of the values of the optical signals. I binned the reconstructed MICH/PRCL in a 51x51 grid, and computed the mean value of all optical signals for each bin. The result is shown in the plot below, directly compared with the expectation from a simulation.

The power signals (POP_DC, AS_DC, PO22_Q) looks reasonably good. REFL11_I/Q also looks good (please note that due to an early mistake in my code, I reversed the convention for I/Q, so PRCL signal is maximized in Q instead than in I). The 55MHz signals look a bit less clear...

Steps forward

  • I'm quite confident in the tuning of demodulation phase and signs for REFL11 and POP22, but less so for REFL55 and not sure at all for AS55. So it would be useful to measure a full sensing matrix of PRCL and MICH against those signals, to compare with my simulation
  • I'm working on an idea to fine tune the DNN using the real interferometer data, more to follow when the idea crystallizes in a clear form.
  13273   Wed Aug 30 10:54:26 2017 gautamUpdateCDSslow machine bootfest

MC autolocker and FSS loops were stuck because c1psl was unresponsive. I rebooted it and did a burtrestore to enable PSL locking. Then the IMC locked fine.

c1susaux and c1iscaux were also unresponsive so I keyed those crates as well, after taking the usual steps to avoid ITMX getting stuck - but it still got stuck when the Sat. Box. connectors were reconnected after the reboot, so I had to shake it loose with bias slider jiggling. This is annoying and also not very robust. I am afraid we are going to knock the ITMX magnets off at some point. Is this problem indicative of the fact that the ITMX magnets were somehow glued on in a skewed way? Or can we make the situation better by just tweaking the OSEM-holding fixtures on the cage?

In any case, I've started listing stuff down here for things we may want to do when we vent next.

 

  13272   Wed Aug 30 06:45:32 2017 KevinSummaryPEMNew Heater Circuit

I changed the heater circuit described in this elog to a current sink. The new and old circuits are shown in the attachment. The heater is R_h and is currently 24Ω; the sense resistor R is currently 6Ω. The op-amp is still an OP27 and the MOSFET is still an IRF630.

The current through the old circuit was saturating because the gate voltage on the MOSFET was saturating at the op-amp supply rails. This is because the source voltage is relatively high: V_S = I(R + R_h).

In the new circuit the source voltage is lower and the op-amp can thus drive a large enough V_{GS} to draw more current (until the power supply saturates at 25V/30Ω = 0.8A in this case). The source and DAC voltages are equal in this caseV_{\mathrm{DAC}} = V_S and so the current is I = V_{\mathrm{DAC}}/R. Since this is the same current through the heater, the drain voltage is V_D = V_{cc} - IR_h. I observed this behavior in this circuit until the power supply saturated at 0.8A. Note that when this happens V_D = V_S and the gate voltage saturates at the supply rails in an attempt to supply the necessary current.

Attachment 1: circuit.pdf
circuit.pdf
  13271   Tue Aug 29 21:36:59 2017 johannesUpdatePSLPSL table auxiliary NPRO
Quote:

 there is a large (~1 cm) aperture Pockels cell that Frank Siefert was using for making pulses to damage photo diodes. There is a DEI Pulser unit near the entrance to the QIL in Bridge which can drive it.

I'll look for it tomorrow, but I haven't given up on the AOMs yet. I swapped in the ISOMET modulator today and saw the same behavior, both in 0th and 1st order. The fall time is pretty much identical. Gautam saw no such thing in the PSL AOM using the same photodetector.

1st order diffracted                                                          0th order

     

In the meantime I prepared the fiber mode-matching but realized in the process that I had mixed up some lenses. As a result the beam did not have a waist at the AOM location and thus didn't have the intended size, although I doubt that this would cause the slower decay. I'll fix it tomorrow, along with setting up the fiber injection, beat note with the PSL, and routing the fiber if possible.

  13270   Tue Aug 29 20:04:09 2017 ranaUpdatePSLPSL table auxiliary NPRO

I don't understand why the 1st order diffracted beam doesn't go to zero when you shut off the drive. My guess is that the standing acoustic wave in the AO crystal needs some time to decay: f = 40 MHz, tau = 1 usec... Q ~ 100. Perhaps, the crystal is damped by the PZT and ther output impedance of the mini-circuits switch is different from the AO driver.

In any case, if you need a faster shut off, or want something that more cleanly goes to zero, there is a large (~1 cm) aperture Pockels cell that Frank Siefert was using for making pulses to damage photo diodes. There is a DEI Pulser unit near the entrance to the QIL in Bridge which can drive it.

  13269   Tue Aug 29 15:41:17 2017 KiraSummaryPEMheater circuit

I worked with Kevin and Gautam to create a heater circuit. The first attachment is Kevin's schematic of the circuit. The OP amp connects to the gate of the power MOSFET, and the power supply connects to the drain, while the source goes into the heater. We set the power supply voltage to 22V and varied the voltage of the input to the OP amp. At 6V to the OP amp, we got a current of 0.35A flowing through the heater and resistor. This was the peak current we got due to the OP amp being saturated (an increase in either of the power supplies did not change the current), but when we increased the voltage of the supply rails of the OP amp from 15V to 20V, we got a current of 0.5A. We would want a higher current than this, so we will need to get a different OP amp with a higher max voltage rating, and a resistor that can take more power than this one (it currently takes 5W of power, and is the best one we could find).

Kevin and I created a simulation of this circuit using CircuitLab to understand why the current was so low (second attachment). The horizontal axis is the voltage we supply to the OP amp. The blue line shows the voltage at the point between the output of the OP amp and the gate of the MOSFET. The orange line is the voltage at the point between the source of the MOSFET and the heater. The brown line is the voltage at the point between the heater and resistor. Thus, we can see that saturation occurs at about 2.1V. At that point, the gate-source voltage is the difference between the blue curve and the orange curve, which is about 4V, which is what we measured. Likewise, the voltage across the heater is the difference between the orange curve and the brown curve, which comes out to around 8V, which is also what we measured. Lastly, the voltage across the resistor is the brown curve, which is about 2V, which matches our observations. The circuit works as it should, but saturates too soon to get a high enough current out of it.

Gautam noted that it is important to measure the current correctly. We can't just use an ammeter and place it across the resistor or heater, because the internal resistance of the ammeter (~0.5 ohm) is comparable to the resistance we want to measure, so the current gets split between the circuit and the ammeter and we get an equivalent resistance of 1/R = 1/R0 + 1/Ra, where R0 is the resistance of the part we want to measure the current across, and Ra is the ammeter resistance. Thus, the new resistance will be lower and the ammeter will show a higher current value than what is actually there. So to accurately measure the current, we must place the ammeter in series with the part we want to measure. We initially got a 1A reading on the heater, which was not correct, and our setup did not heat up at all basically. When we placed the ammeter in series with the heater, we got only 0.35A.

The last two images are the setup for testing of the heater. We wrapped it around an aluminum piece and covered it with a few layers of insulating material. We can stick a thermometer in between the insulation and heater to see the temperature change. In later tests, we may insulate the whole piece so that less heat gets dissipated. In addition, we used a heat sink and thermal paste to secure the MOSFET to it, as it got very hot.

Our next steps will be to get a resistor and an OP amp that are better suited for our purposes. We will also run simulations with components that we choose to make sure that it can provide the desired current of 1A (the maximum output of the power supply is 24V, and the heater is 24 ohm, so max current is 1A). Kevin is working on that now.

Attachment 1: heater_circuit.pdf
heater_circuit.pdf
Attachment 2: simulation.png
simulation.png
Attachment 3: heater_setup.jpg
heater_setup.jpg
Attachment 4: IMG_20170829_131126.jpg
IMG_20170829_131126.jpg
  13268   Tue Aug 29 15:35:19 2017 SteveUpdateVACvacuum pump specifications & manuals

RP1 and RP3 roughing pump manual of Leybold D30A oily rotory pump

Fore pump of TP2 & TP3 Varian SH-100 Dry Scroll

TP2 and TP3 small turbo drag pump  Varian 969-9361 

TP2 and TP3 turbo controller Varian 969-9505

TP1 magnetically suspended turbo pump  Osaka TG390MCAB, sn360 and controller TC010M  and note : this  pump  running on 208VAC single phaseIt is not on the UPS !  

                                                             Osaka Maglev Manual        and Osaka Controller Communication Wiring                                                                                                                                

VC1 cryo pump CTI-Cryogenics Cryo Torr 8 sn 8g23925  SAFETY note: compressor single phase 208VAC and the head driver 3 phase 208VAC Compressor and driver have each separate power cord!

Installed at 40m wiki  also

Quote:

The V1 gate valve specs installed at 40m wiki page. VAT model number 10846-UE44-0007 Our main volume pumping goes through this 8" id gate valve V1 to Maglev turbo  or Cryo pump to  VC1

The ion pumps have 6" id gate valves:VAT 10844-UE44-AAY1,  Pneumatic actuator with position indicator and double acting solenoid valve 115V 60Hz  Purchased 1999 Dec 22

UHV gate valves 2.5" id. VAT  10836-UE44    Pneumatic actuator with position indicator and double acting solenoid valve 115V 60 hz, IFO to RGA  VM1 &  RGA to Maglev  VM2

mini UHV gate valve 1.5" id.   VAT  01032-UE01      2016 cataloge page 14,   manual - no position indicator, VM4  next to manual adjustable fine leak valve to RGA

UHV angle valve 1.5" id, model VAT 28432-GE41, Viton plate seal, pneumatic actuator with position indicator & solenoid valve 115V & single acting closing spring  MEDM screen: VM3,VC2, V3,V4,V5,V6,VA6,V7 & annuloses Each chamber annulos has 2 valves.

UHV angle valve 1.5" id, model VAT 57132-GE05   go page 208,   Metal tip seal, manual actuating only with position indicator,   MEDM screen: roughing RV1 and venting VV1 hand wheel needed to close to torque spec

UHV angle valve  1.5" id. model VAT 28432-GE01           Viton plate seal, manual operation only at IT gauges Hornet & Super Bee and  ion pumps roughing  ports. These are not labeled.

                                                      

The Cryo pump interlock wiring was added too

Note: all moving valve plate seals are single.

 

  13267   Tue Aug 29 15:04:59 2017 gautamUpdateSUSETMY Oplev PIT loop gain changed

Last night, while we were working on the ALS, I noticed the GTRY spot moving around (in PITCH) on the CCD monitor in the control room at ~1-2Hz. The operating condition was that the arm was locked to the IR, and the PSL green shutter was closed, so that only the arm transmissions were visible on the CCD screens. There was no such noticable movement of the GTRX spot. When looking at the out-of-loop ALS nosie in this configuration (but now with the PSL green shutter open of course), the Y arm ALS noise at low frequencies was much worse than the X arm.

Today, I looked into this a little more. I first checked that the Y-endtable enclosure was closed off as usual (as I had done some tweaking to the green input pointing some days ago). There are various green ghost beams on the Y-endtable. When time permits, we should make an effort to cleanly dump these. But the enclosure was closed as usual.

Then I looked at the in-loop Oplev error signal spectra for the ITMY and ETMY Oplev loops. There was high coherence between ETMYP Oplev error signal and GTRY. So I took a loop transfer function measurement - the upper UGF was around 3.5Hz. I increased the loop gain such that the upper UGF was around 4.5Hz, with phase margin ~30degrees. Doing so visibly reduced the angular movement of the GTRY spot on the CCD. Attachment #1 shows the Oplev loop TF after the gain increase, while Attachment #2 compares the GTRX and GTRY spectra (DC value is approximately the same for both, around 0.4). GTRY still seems a bit noisier at low frequencies, but the out-of-loop ALS noise for the Y arm now lines up much more closely with its reference trace from a known good time. 

Quote:
 

Y-arm ALS wasn't so stellar tonight, especially at low frequencies. I can see the GTRY spot moving on the CCD monitor, so something is wonky. To be investigated.

 

Attachment 1: ETMY_OLPIT.pdf
ETMY_OLPIT.pdf
Attachment 2: GTR_comparison.pdf
GTR_comparison.pdf
  13266   Tue Aug 29 02:08:39 2017 gautamUpdateALSFiber ALS noise measurement

I was having a chat with EricQ about this today, just noting some points from our discussion down here so that I remember to look into this tomorrow.

  • I believe that currently, the channels C1:ALS-BEATX_FINE_PHASE_OUT_HZ_DQ and the Y arm analog read out the frequency of the green beat, in Hz.
  • In the comparison I plotted, I WRONGLY divided the spectrum of the IR beat by 2, instead of multiplying in by 2, which is what should actually be done for an apples-to-apples comparison.
  • The deeper question is, what should this channel actually readout?
  • Looking at my codes from past arm scans etc, I see that I am dividing the downloaded data by 2 in order to convert the X-axis of these scans to "IR Hz". But this should really be all we care about.
  • So I think I will have to re-do the cts-to-Hz calibration in the ALS models. It should be possible to do ~10FSR scans with the IR beat, and then we can use the sideband resonances (presumably the sideband frequencies are known with better precision than the arm length, and hence the FSR) to calibrate the phase tracker.
  • I don't think this changes the fact that the Fiber ALS situation has been improved - but I will have to repeat the measurement to be sure. The improvement may not be as stellar as I tried to sell in my previous elog sad.

    Other thoughts: 

  • Can we make use of the Jetstor raid array for some kind of consolidated 40m CDS backup system? Once we've gotten everything of interest out of it...

  13265   Tue Aug 29 01:52:22 2017 gautamUpdateSUSTest mass actuator calibration

[ericq, gautam]

Tonight, we decided to double-check the POX counts-to-meters conversion.

It is unclear when this was last done, and since I modified the coil driver electronics for the ITMs and BS recently, I figured it would be useful to get this calibration done. The primary motivation was to see if we could resolve the discrepancy between the current ALS noise (using POX as a sensor) compared to the Izumi et. al. plot.

Because we are planning to change the coil driver electronics further soon anyways, we decided to do the calibration at a single frequency for tonight. For future reference, the extension of this method to calibrate the actuator over a wider range of frequencies is here. The procedure followed, and the relevant numbers from tonight, are as follows.

Procedure:

  1. Set dark offsets on all DCPDs and LSC PDs.
  2. Look at the free swinging Michelson signal on ASDC.
    • For tonights test, ASDC was derived from the AS55 photodiode.
    • The AS110 photodiode actually has more light on it, but we think that the ADC that the DCPD board is interfaced to is running on 0-2V rather than 0-10V, as the signal seemed to saturate around 2000 counts. It is unclear whether the actual photodiode is saturating, to be investigated.
    • So we decided to use ASDC from AS55 photodiode with 15dB whitening gain.
    • There is also some issue with the whitening filter (not whitening gain) on ASDC - engaging the whitening shifts the DC offset. This has to be investigated while we get stuck into the LSC electronics.
  3. Look at the peak-to-peak swing of ASDC. Use algebraic expression for reflected power from Michelson interferometer to calibrate the ASDC slope at Michelson half-fringe. For the test tonight, ASDC_max = 1026 counts, ASDC_min = 2 counts.
  4. Lock the Michelson at half-fringe, with ASDC as the error signal.
    • Zero out the MICH elements in the RFPD input matrix.
    • Set the matrix element from ASDC to MICH in the DCPD LSC input matrix to 1.
    • The servo gain used was +0.005 on the MICH_A servo path.
    • A low-frequency boost was turned on.
  5. Use the sensing matrix infrastructure to drive a line in the optic of interest.
    • Tonight, we looked at ITMX and ITMY.
    • The line was driven at 311.1Hz, and the amplitude was 300 counts.
    • Download 60secs of ASDC data, demodulate at the driven frequency to find the peak height in counts, and using the slope of ASDC (in cts/m) at the Michelson half-fringe, calculate the actuator gain in m/cts.
    • ITMY: 2.55e-9 / f^2 m/count
    • ITMX: 2.65e-9 / f^2 m/count
    • These numbers kind of make sense - the previous numbers were ~5nm/f^2 /ct, but I removed an analog gain of x3 in this path. Presumably there has been some change in the N/A conversion factor - perhaps because of a change in the interaction between the optics' face magnets and the static magnetic field in the OSEMs?
  6. Lock the arms with POX/POY, and drive the newly calibrated ITMs.
    • So we know how many meters we are driving the ITMs by.
    • Looking at POX/POY, we can calibrate these into meters/count.
    • Both POX and POY were whitened.
    • POX whitening gain = +30dB, POY whitening gain = +18dB.
    • ITMX and ITMY were driven at 311.1Hz, with amplitude = 2counts.
    • Download 60 secs of data, demodulate at the drive frequency to find the peak height, and use the known ITM actuator gains to calibrate POX and POY.
    • POX: 7.34e-13 m / count (approx. 5 times less than the number in the Foton filter bank in the C1:CAL-CINV model).
    • POY: 1.325e-13 m / count
    • We did not optimize the demod phases for POX/POY tonight. 

Once these calibrations were updated, we decided to control the arms with ALS, and look at the POX spectrum. Y-arm ALS wasn't so stellar tonight, especially at low frequencies. I can see the GTRY spot moving on the CCD monitor, so something is wonky. To be investigated. But the X arm ALS noise looked pretty good.

Seems like updating the calibration did the job; see the attached comparison plot.

Attachment 1: ALS_comparison.pdf
ALS_comparison.pdf
  13264   Mon Aug 28 23:22:56 2017 johannesUpdatePSLPSL table auxiliary NPRO

I moved the axuiliary NPRO to the PSL table today and started setting up the optics.

The Faraday Isolator was showing a pretty unclean mode at the output so I took the polarizers off to take a look through them, and found that the front polarizer is either out of place or damaged (there is a straight edge visible right in the middle of the aperture, but the way the polarizer is packaged prevents me from inspecting it closer). I proceeded without it but left space so an FI can be added in the future. The same goes for the broadband EOM.

There are two spare AOMs (ISOMET and Intraaction, both resonant at 40MHz) available before we have to resort to the one currently installed in the PSL.

I installed the Intraaction AOM first and looked at the switching speed of its first order diffracted beam using both its commercial driver and a combination of minicircuits components. Both show similar behavior. The fall time of the initial step is ~110ns in both cases, but it doesn't decay rapidly no light but a slower exponential. Need to check the 0 order beam and also the other AOM.

Intraaction Driver

   

Mini Circuits Drive

   

  13263   Mon Aug 28 17:13:57 2017 ericqUpdateCDS40m files backup situation

In addition to bootable full disk backups, it would be wise to make sure the important service configuration files from each machine are version controlled in the 40m SVN. Things like apache files on nodus, martian hosts and DHCP files on chiara, nds2 configuration and init scripts on megatron, etc. This can make future OS/hardware upgrades easier too.

  13262   Mon Aug 28 16:20:00 2017 gautamUpdateCDS40m files backup situation

This elog is meant to summarize the current backup situation of critical 40m files.

What are the critical filesystems? I've also indicated the size of these disks and the volume currently used, and the current backup situation. 

Name

Disk Usage

Description / remarks

Current backup status

FB1 root filesystem 1.7TB / 2TB
  • FB1 is the machine that hosts the diskless root for the front end machines
  • Additionally, it runs the daqd processes which write data from realtime models into frame files
Not backed up
/frames up to 24TB
  • This is where the frame files are written to 
  • Need to setup a wiper script that periodically clears older data so that the disk doesn't overflow.

Not backed up 

LDAS pulls files from nodus daily via rsync, so there's no cron job for us to manage. We just allow incoming rsync.

Shared user area 1.6TB / 2TB
  • /home/cds on chiara
  • This is exported over NFS to 40m workstations, FB1 etc.
  • Contains user directories, scripts, realtime models etc.

Local backup on /media/40mBackup on chiara via daily cronjob

Remote backup to ldas-cit.ligo.caltech.edu::40m/cvs via daily cronjob on nodus

Chiara root filesystem 11GB / 440GB
  • This is the root filesystem for chiara
  • Contains nameserver stuff for the martian network, responsible for rsyncing /home/cds
Not backed up
Megatron root filesystem 39GB / 130GB
  • Boot disk for megatron, which is our scripts machine
  • Runs MC autolocker, FSS loops etc.
  • Also is the nds server for facilitating data access from outside the martian network
Not backed up
Nodus root filesystem 77GB / 355GB
  • This is the boot disk for our gateway machine
  • Hosts Elog, svn, wikis
  • Supposed to be responsible for sending email alerts for NFS disk usage and vacuum system N2 pressure
Not backed up
JETSTOR RAID Array 12TB / 13TB
  • Old /frames
  • Archived frames from DRFPMI locks
  • Long term trends

Currently mounted on Megatron, not backed up.

Then there is Optimus, but I don't think there is anything critical on it. 

So, based on my understanding, we need to back up a whole bunch of stuff, particularly the boot disks and root filesystems for Chiara, Megatron and Nodus. We should also test that the backups we make are useful (i.e. we can recover current operating state in the event of a disk failure).

Please edit this elog if I have made a mistake. I also don't have any idea about whether there is any sort of backup for the slow computing system code.

  13261   Mon Aug 28 10:51:21 2017 steveUpdateSUSETMX damping recovered
Attachment 1: ETMX_restored.png
ETMX_restored.png
  13260   Mon Aug 28 10:28:07 2017 SteveUpdateVACRGA scan 20 day after 17 Torr

The RGA was turned on 7 days ago.  It's 46 C now.   The X-arm room tem ~20 C

IFO pressure 6.5e-6 Torr at IT-Hornet gauge. Valve configuration vacuum normal.

Attachment 1: 20d_after_17_torr.png
20d_after_17_torr.png
  13258   Mon Aug 28 08:47:32 2017 JamieSummaryLSCFirst cavity length reconstruction with a neural network
Quote:

Phenomenal!

truly.

  13257   Sun Aug 27 11:57:31 2017 ranaUpdateALSFiber ALS noise measurement

It seems like the main contribution to the RMS comes from the high frequency bump. When using the ALS loop to lock the arm to the beat, only the stuff below ~100 Hz will matter. Interesting to see what that noise budget will show. Perhaps the discrepancy between inloop and out of loop will go down.

  13256   Sat Aug 26 09:56:34 2017 GabrieleSummaryLSCFirst cavity length reconstruction with a neural network

Update

I included the 55 MHz sideband and higher order modes in my training examples. To keep things simple, I just assumed there are higher order modes up to n+m=4 in the input beam. The power in each HOM is randomly chosen from a random gaussian distribution with width determined from experimental cavity scans. I used a value of 0.913+-0.01 rad for the Gouy phase (again estimated from cavity scans, but in reasonable agreement with the nominal radius of curvature of ETMX)

Results are improved. The plot belows show the performance of the neural network on 100s of experimental data

For reference, the plots below show the performance of the same network on simulated data (that includes sensing noise but no higher order modes)

  13255   Fri Aug 25 17:11:07 2017 ranaUpdateALSFiber ALS noise measurement

Is it better to mount the box in the PSL under the existing shelf, or in a nearby PSL rack?

Quote:

 

Further characterization needs to be done, but the results of this test are encouraging. If we are able to get this kind of out of loop ALS noise with the IR beat, perhaps we can avoid having to frequently fine-tune the green beat alignment on the PSL table. It would also be ideal to mount this whole 1U setup in an electronics rack instead of leaving it on the PSL table

 

  13254   Fri Aug 25 15:54:14 2017 gautamUpdateALSFiber ALS noise measurement

[Kira, gautam]

Attachment #1 - Photo of the revamped beat setup. The top panel has to be installed. New features include:

  • Regulated power supply via D1000217.
  • Single power switch for both PDs.
  • Power indicator LED.
  • Chassis ground isolated from all other electronic grounds. For this purpose, I installed all the elctronics on a metal plate which is only connected to the chassis via nylon screws. The TO220 package power regulator ICs have been mounted with the TO220 mounting kits that provide a thin piece of plastic that electrically insulates its ground from the chassis ground.
  • PD outputs routed through 20dB coupler on front panel for diagnostic purposes.
  • Fiber routing has been cleaned up a little. I installed a winding fixture I got from Johannes, but perhaps we can install another one of these on top of the existing one to neaten up the fiber layout further.
  • 90-10 light splitter (meant for diagnostic purposes) has been removed because of space constraints. 

Attachment #2 - Power budget inside the box. Some of these FC/APC connectors seem to not offer good coupling between the two fibers. Specifically, the one on the front panel meant to accept the PSL light input fiber seems particularly bad. Right now, the PSL light is entering the box through one of the front panel connectors marked "PSL + X out". I've also indicated the beat amplitude measured with an RF analyzer. Need to do the math now to confirm if these match the expected amplitudes based on the power levels measured.

Attachment #3 - We repeated the measurement detailed here. The X arm (locked to IR) was used for this test. The "X" delay line electronics were connected to the X green beat PD, while the "Y" delay line electronics were connected to the X IR beat PD. I divided the phase tracker Hz calibration factor by 2 to get IR Hz for the Y arm channels. IR beat was at ~38MHz, green beat was at ~76MHz. The broadband excess noise seen in the previous test is no longer present. Indeed, below ~20Hz, the IR beat seems less noisy. So seems like the cleaning / electronics revamp did some good. 

Further characterization needs to be done, but the results of this test are encouraging. If we are able to get this kind of out of loop ALS noise with the IR beat, perhaps we can avoid having to frequently fine-tune the green beat alignment on the PSL table. It would also be ideal to mount this whole 1U setup in an electronics rack instead of leaving it on the PSL table.

Quote:

Photos + power budget + plan of action for using this box to characterize the green PDH locking to follow. 

GV Edit: I've added better photos to the 40m Google Photos page. I've also started a wiki page for this box / the proposed IR ALS  system. For the moment, all that is there is the datasheet to the Fiber Couplers used, I will populate this more as I further characterize the setup.

Attachment 1: IMG_7497.JPG
IMG_7497.JPG
Attachment 2: FOL_schematic.pdf
FOL_schematic.pdf
Attachment 3: 20170825_IR_ALS.pdf
20170825_IR_ALS.pdf
  13253   Fri Aug 25 11:11:26 2017 gautamUpdateGeneralMC1 kicked again

Looks like MC1 got another big kick just under 4 hours ago. None of the other optics show any evidence of a glitch so it seems unlikely that this was some sort of global event. It's been well behaved for ~2weeks now. IMC was unlocked. I manually re-aligned MC1, at which point the autolocker was able to lock the IMC.

Looking at this plot, it seems that LR and UL coils seem to have the largest kicks. UR barely saw it. Not sure what (if anything) to make of this - apparently the optic moved by ~20urad with the UR magnet approximately the pivot.

Attachment 1: MC1_glitch.png
MC1_glitch.png
  13252   Fri Aug 25 01:20:52 2017 gautamUpdateLSCDRMI locking attempt

I tried some DRMI locking again tonight, but had no success. Here is the story.

  • I started out by going to the AS table and measuring the light level on the REFL55 photodiode (with PRM aligned and the PRC flashing, but LSC disabled).
    • The Ophir power meter reads 13mW
    • The DC output of the photodiode shows ~500mV on an oscilloscope.
    • Both of these numbers line up well with measurements I made in April/May.
  • Returned to the control room and aligned the IFO for DRMI locking - but LSC servos remained disabled.
    • At the nominal REFL55 whitening level of +18dB, the REFL 55 signals saturated the ADC (confirmed by looking at the traces on dataviewer).
    • But the signals still looked like PDH error signals.
    • Lowering the whitening gain to 6dB makes the PDH error signal horns peak around 20,000 counts.
    • Could this be indicative of problems with either the analog whitening gain switching or the LSC Demod Boards? To be investigated.
  • Tried enabling LSC servos with same settings with which I had success right up till a couple of months ago, but had no success.
    • If it is true that the REFL55 signal is getting amplified because of some gain stage not being switched correctly, I should still have been able to lock the SRC with a lowered loop gain - but even lowering the gain by a factor of 10 had no effect on the locking success rate.

Looks like I will have to embark on the REFL55 LSC electronics investigation. I was able to successfully lock the PRC on carrier and sideband, and the Michelson lock also seems to work fine, all of which seem to point to a hardware problem with the REFL55 signal chain.

I did a quick check by switching the output of the REFL55 demod board to the inputs normally used by AS55 signals on the whitening board. Setting the whitening gain to +18dB for these channels had the same effect - ADC overflow galore. So looks like the whitening board isn't to blame. I will have to check the demod board out.

 

  13251   Thu Aug 24 18:51:57 2017 KojiSummaryLSCFirst cavity length reconstruction with a neural network

Phenomenal!

  13250   Thu Aug 24 18:02:16 2017 GabrieleSummaryLSCFirst cavity length reconstruction with a neural network

1) Introduction

In brief, I trained a deep neural network (DNN) to recosntuct the cavity length, using as input only the transmitted power and the reflection PDH signals. The training was performed with simulated data, computed along 0.25s long trajectories sampled at 8kHz, with random ending point in the [-lambda/4, lambda/4] unique region and with random velocity.

The goal of thsi work is to validate the whole approach of length reconstruction witn DNN in the Fabry-Perot case, by comparing the DNN reconstruction with the ALS caivity lenght measurement. The final target is to deploy a system to lock PRMI and DRMI. Actually, the Fabry-Perot cavity problem is harder for a DNN: the cavity linewidth is quite narrow, forcing me to use very high sampling frequency (8kHz) to be able to capture a few samples at each resonance crossing. I'm using a recurrent neural network (RNN), in the input layers of the DNN, and this is traine using truncated backpropagation in time (TBPT): during training each layer of RNN is unrolled into as many copies as there are input time samples (8192 * 0.25 = 2048). So in practice I'm training a DNN with >2000 layers! The limit here is computational, mostly the GPU memory. That's why I'm not able to use longer data stretches.

But in brief, the DNN reconstruction is performing well for the first attempt.

2) Training simulation

In the results shown below, I'm using a pre-trained network with parameters that do not match very well the actual data, in particular for the distribution of mirror velocity and the sensing noises. I'm working on improving the training.

I used the following parameters for the Fabry-Perot cavity:

The uncertaint is assumed to be the 90% confidence level of a gaussian distribution. The DNN is trained on 100000 examples, each one a 0.25/8kHz long trajectory with random velocity between 0.1 and 5 um/s, and ending point distributed as follow: 33% uniform on the [-lambda/4, lambda/4] region, plus 33% gaussian distribution peaked at the center with 5 nm width. In addition there are 33% more static examples, distributed near the center. 

For each point along the trajectory, the signals TRA, POX11_I and POX11_Q are computed and used as input to the DNN.

3) Experimental data

Gautam collected about 10 minutes of data with the free swinging cavity, with ALS locked on the arm. Some more data were collected with the cavity driven, to increase the motion. I used the driven dataset in the analysis below.

3.1) ALS calibration

The ALS signal is calibrated in green Hz. After converting it to meters, I checked the calibration by measuring the distance between carrier peaks. It turned out that the ALS signal is undercalibrated by about 26%. After correcting for this, I found that there is a small non-linearity in the ALS response over multiple FSR. So I binned the ALS signal over the entire range and averaged the TRA power in each bin, to get the transmission signals as a function of ALS (in nm) below:

I used a peak detection algorithm to extract the carrier and 11 MHz sideband peaks, and compared them with the nominal positions. The difference between the expected and measured peak positions as a function of the ALS signal is shown below, with a quadratic fit that I used to improve the ALS calibration

The result is

z_initial = 1e9 * L*lamba/c *1.26. * ALS
z_corrected = 2.1e-06 z^2  -1.9e-02 z  -6.91e+02

The ALS calibrated z error from the peak position is of the order of 3 nm (one sigma)

3.2) Mirror velocity

Using the calibrated ALS signal, I computed the cavity length velocity. The histogram below shows that this is well described by a gaussian with width of about 3 um/s. In my DNN training I used a different velocity distribution, but this shouldn't have a big impact. I'm retraining with a different distirbution.

4) DNN results

The plot below shows a stretch of time domain DNN reconstruction, compared with the ALS calibrated signal. The DNN output is limited in the [-lambda/4, lambda/4] region, so the ALS signal is also wrapped in the same region. In general the DNN reconstruction follows reasonably well the real motion, mostly failing when the velocity is small and the cavity is simultanously out of resonance. This is a limitation that i see also in simulation, and it is due to the short training time of 0.25s.

I did not hand-pick a good period, this is representative of the average performance. To get a better understanding of the performance, here's a histogram of the error for 100 seconds of data:

The central peak was fitted with a gaussian, just to give a rough idea of its width, although the tails are much wider. A more interesting plot is the hisrogram below of the reconstructed position as a function of the ALS position, Ideally one would expect a perfect diagonal. The result isn't too far from the expectation:

The largest off diagonal peak is at (-27, 125) and marked with the red cross. Its origin is more clear in the plot below, which shows the mean, RMS and maximum error as a function of the cavity length. The second peak corresponds to where the 55 MHz sideband resonate. In my training model, there were no 55 MHz sidebands nor higher order modes. 

5) Conclusions and next steps

The DNN reconstruction performance is already quite good, considering that the DNN couldn't be trained optimally because of computation power limitations. This is a validation of the whole idea of training the DNN offline on a simulation and then deploy the system online.

I'm working to improve the results by

  • training on a more realistic distribution of velocity
  • adding the 55 MHz sidebands
  • maybe adding HOMs
  • tune the DNN architecture

However I won't spend too much time on this, since I think the idea has been already validated.

 

  13249   Thu Aug 24 17:36:11 2017 gautamUpdateCDSFSS Slow Python maintenance

A couple of weeks ago, I was trying to modernize the python version of the FSS Slow temperature control loops, when I accidentally ended up deleting it frown. There was no svn backup. So the old Perl PID script has been running for the last few days.

Today, I checked out the latest version that Andrew and co. have running in the PSL lab. I had to make some important modifications for the script to work for the 40m setup.

  1. The script is conveniently setup in a way that the channels it needs to read from / write to are read in from an .ini file. I renamed all the channels to match the appropriate 40m ones.
  2. We don't have a soft epics channel in which to define the setpoint for our PID servo (which is 0). Rather than poke around with slow machine EPICS records, I simply commented out this line in the script and included the hard-coded value of 0. When we modernize to the Acromag era, we can setup an EPICS channel + MEDM slider for the setpoint.
  3. The way the Perl script was setup, the error signal was pre-scaled by a factor of 0.01, supposedly to make the PID gains be of order 1. For consistency, I re-inserted this scaling, which awade and co. had removed.
  4. Modified the FSSslowPy.init file to call the script in accordance with the new syntax:
python FSSSlow.py -i FSSSlowPy.ini

Then I stopped the Perl process on megatron by running

sudo initctl stop FSSslow

and started the Python process by running

sudo initctl start FSSslowPy

I have now committed the files FSSSlow.py and FSSSlowPy.ini to the 40m svn.  Things seem to be stable for the last 20 mins or so, let's keep an eye on this though - although we had been running the Python PID loop for some months, this version is a slightly modified one. 

The initctl stuff still isn't very robust - I think both the Autolocker and the FSS slow servos have to be manually restarted if megatron is shutdown/restarted for whatever reason. It doesn't seem to be a problem with the initctl routine itself - looking at the logs, I can see that init is trying to start both processes, but is failing to do so each time. To be investigated. The wiki procedure to restart this process is up to date.

GV Edit 0000 25 Aug 2017: I had to add a line to the script that checks MC transmission before enabling the PID loop. Change has been committed to svn. Now, when the MC loses lock or if the PSL shutter is kept closed for an extended period of time, the temperature loop doesn't rail.

  13248   Thu Aug 24 00:39:47 2017 gautamUpdateLSCDRMI locking attempt

Since the single arm locking and dither alignment seemed to work alright after the CDS overhaul, I decided to try some recycling cavity locking tonight.

  • First, I locked single arms, ran dither alignment servos, and centered all test mass Oplevs. Note: the X arm dither alignment doesn't seem to work if we use the High-Gain Thorlabs PD as the Transmission PD. The BS loops just seem to pick up large offsets and the alignment actually degrades over a couple of minutes. This needs to be investigated.
  • Next, to get good PRM alignment, I manually moved the EPICS sliders till the REFL spot became roughly centered on the CCD screen.
  • Then I tried locking PRMI on carrier using the usual C1IFOConfigure script - the lock was caught within ~30 seconds.
  • The PRCL and MICH dither servo scripts also ran fine.
    • Centered PRM Oplev.
  • Next, I tried enabling the PRC angular feedforward.
    • OAF model does not automatically revert to its safe.snap configuration on model reboot, so I first manually did this such that the correct filter banks were enabled.
    • I was able to turn on the angular feedforward without disturbing the PRMI carrier lock. The angular motion of the POP spot on the CCD monitor was visibly reduced.
  • At this point I decided to try DRMI locking.
    • I centered the beam on the AS PDs with the simple Michelson.
    • Centered the beam on the REFL PDs with PRM aligned and PRC flashing through resonances.
    • Restored SRM alignment by eye with EPICS sliders.
    • Cavity alignment seemed alright - so I tried to lock DRMI with the old settings (i.e. from DRMI 1f locking a couple of months ago). But I had no success.
    • The behaviour of REFL55 (used for SRCL control) has changed dramatically - the analog whitening gain for this PD used to be +18dB, but at this setting, there are frequent ADC overflows. I had to reduce the whitening gain to +6dB to stop the ADC overflows. I also checked to make sure that the whitening setting was "manual" and not triggered.

Why should this have changed? I was just on the AS table and did re-center the beam onto the REFL 55 RFPD, but I had also done this in April/May when I was last doing DRMI locking. But I can't explain the apparent factor of ~4 increase in light level. I think I have some measurements of the light levels at various PDs from April 2017, I will see how the present levels line up.

Of course dataviever won't cooperate when I am trying to monitor testpoints.

I may be missing something obvious, but I am quitting for tonight, will look into this more tomorrow.


Unrelated to this work: looking at the GTRY spot on the CCD monitor, there seems to be some excess angular motion. Not sure where this is coming from. In the past, this sort of problem has been symptomatic of something going wonky with the Oplev loops. But I took loop measurements for ITMY and ETMY PIT and YAW, they look normal. I will investigate further when I am doing some more ALS work.

  13246   Wed Aug 23 17:22:36 2017 gautamUpdateALSFiber ALS - reinstalled

I completed the revamp of the box, and re-installed the box on the PSL table today. I think it would be ideal to install this on one of the electronic racks, perhaps 1X2 would be best. We would have to re-route the fibers from the PSL table to 1X2, but I think they have sufficient length, and this way, the whole arrangement is much cleaner.

Did a quick check to make sure I could see beat notes for both arms. I will now attempt to measure the ALS noise with this revamped box, to see if the improved power supply and grounding arrangement, as well as fiber cleaning, has had any effect.

Photos + power budget + plan of action for using this box to characterize the green PDH locking to follow. 

For quick reference: here is the AM/PM measurement done when we re-installed the repaired Innolight NPRO on the new X endtable.

  13245   Wed Aug 23 10:11:46 2017 SteveUpdateVACvacuum valve specifications

The V1 gate valve specs installed at 40m wiki page. VAT model number 10846-UE44-0007 Our main volume pumping goes through this 8" id gate valve V1 to Maglev turbo  or Cryo pump to  VC1

The ion pumps have 6" id gate valves:VAT 10844-UE44-AAY1,  Pneumatic actuator with position indicator and double acting solenoid valve 115V 60Hz  Purchased 1999 Dec 22

UHV gate valves 2.5" id. VAT  10836-UE44    Pneumatic actuator with position indicator and double acting solenoid valve 115V 60 hz, IFO to RGA  VM1 &  RGA to Maglev  VM2

mini UHV gate valve 1.5" id.   VAT  01032-UE01      2016 cataloge page 14,   manual - no position indicator, VM4  next to manual adjustable fine leak valve to RGA

UHV angle valve 1.5" id, model VAT 28432-GE41, Viton plate seal, pneumatic actuator with position indicator & solenoid valve 115V & single acting closing spring  MEDM screen: VM3,VC2, V3,V4,V5,V6,VA6,V7 & annuloses Each chamber annulos has 2 valves.

UHV angle valve 1.5" id, model VAT 57132-GE05   go page 208,   Metal tip seal, manual actuating only with position indicator,   MEDM screen: roughing RV1 and venting VV1 hand wheel needed to close to torque spec

UHV angle valve  1.5" id. model VAT 28432-GE01           Viton plate seal, manual operation only at IT gauges Hornet & Super Bee and  ion pumps roughing  ports. These are not labeled.

                                                      

The Cryo pump interlock wiring was added too

Note: all moving valve plate seals are single.

  13244   Tue Aug 22 23:27:14 2017 ranaUpdateALSALS OLTFs

Didn't someone look at what the OLG req. should be for these servos at some point? I wonder if we can make a parallel digital path that we switch on after green lock. Then we could make this a simple 1/f box and just add in the digital path (take analog control signal into ADC, filter, and then sum into the control point further down the path to the laser) for the low frequency boost.

  13243   Tue Aug 22 18:36:46 2017 gautamUpdateComputersAll FE models compiled against RCG3.4

After getting the go ahead from Jamie, I recompiled all the FE models against the same version of RCG that we tested on the c1iscex models.

To do so:

  • I did rtcds make and rtcds install for all the models.
  • Then I ssh-ed into the FEs and did rtcds stop all, followed by rtcds start <model> in the order they are listed on the CDS overview MEDM screen (top to bottom).
  • During the compilation process (i.e. rtcds make), for some of the models, I got some compilation warnings. I believe these are related to models that have custom C code blocks in them. Jamie tells me that it is okay to ignore these warnings at that they will be fixed at some point.
  • c1lsc FE crashed when I ran rtcds stop all - had to go and do a manual reboot.
  • Doing so took down the models on c1sus and c1ioo that were running - but these FEs themselves did not have to be robooted.
  • Once c1lsc came back up, I restarted all the models on the vertex FEs. They all came back online fine.
  • Then I ssh-ed into FB1, and restarted the daqd processes - but c1lsc and c1ioo CDS indicators were still red.
  • Looks like the mx_stream processes weren't started automatically on these two machines. Reasons unknown. Earlier today, the same was observed for c1iscey.
  • I manually restarted the mx_stream processes, at which point all CDS indicator lights became green (see Attachment #1).

IFO alignment needs to be redone, but at least we now have a (admittedly rounabout way) of getting testpoints. Did a quick check for "nan-s" on the ASC screen, saw none. So I am re-enabling watchdogs for all optics.

GV 23 August 9am: Last night, I re-aligned the TMs for single arm locks. Before the model restarts, I had saved the good alignment on the EPICs sliders, but the gain of x3 on the coil driver filter banks have to be manually turned on at the moment (i.e. the safe.snap file has them off). ALS noise looked good for both arms, so just for fun, I tried transitioning control of both arms to ALS (in the CARM/DARM basis as we do when we lock DRFPMI, using the Transition_IR_ALS.py script), and was successful.

Quote:

[jamie, gautam]

We tried to implement the fix that Rolf suggested in order to solve (perhaps among other things) the inability of some utilities like dataviewer to open testpoints. The problem isn't wholly solved yet - we can access actual testpoint data (not just zeros, as was the case) using DTT, and if DTT is used to open a testpoint first, then dataviewer, but DV itself can't seem to open testpoints.

Here is what was done (Jamie will correct me if I am mistaken).

  1. Jamie checked out branch 3.4 of the RCG from the SVN.
  2. Jamie recompiled all the models on c1iscex against this version of RCG.
  3. I shutdown ETMX watchdog, then ran rtcds stop all on c1iscex to stop all the models, and then restarted them using rtcds start <model> in the order c1x01, c1scx and c1asx. 
  4. Models came back up cleanly. I then restarted the daqd_dc process on FB1. At this point all indicators on the CDS overview screen were green.
  5. Tried getting testpoint data with DTT and DV for ETMX Oplev Pitch and Yaw IN1 testpoints. Conclusion as above.

So while we are in a better state now, the problem isn't fully solved. 

Comment: seems like there is an in-built timeout for testpoints opened with DTT - if the measurement is inactive for some time (unsure how much exactly but something like 5mins), the testpoint is automatically closed.

 

Attachment 1: CDS_Aug22.png
CDS_Aug22.png
  13242   Tue Aug 22 17:11:15 2017 gautamUpdateComputersc1iscex model restarts

[jamie, gautam]

We tried to implement the fix that Rolf suggested in order to solve (perhaps among other things) the inability of some utilities like dataviewer to open testpoints. The problem isn't wholly solved yet - we can access actual testpoint data (not just zeros, as was the case) using DTT, and if DTT is used to open a testpoint first, then dataviewer, but DV itself can't seem to open testpoints.

Here is what was done (Jamie will correct me if I am mistaken).

  1. Jamie checked out branch 3.4 of the RCG from the SVN.
  2. Jamie recompiled all the models on c1iscex against this version of RCG.
  3. I shutdown ETMX watchdog, then ran rtcds stop all on c1iscex to stop all the models, and then restarted them using rtcds start <model> in the order c1x01, c1scx and c1asx. 
  4. Models came back up cleanly. I then restarted the daqd_dc process on FB1. At this point all indicators on the CDS overview screen were green.
  5. Tried getting testpoint data with DTT and DV for ETMX Oplev Pitch and Yaw IN1 testpoints. Conclusion as above.

So while we are in a better state now, the problem isn't fully solved. 

Comment: seems like there is an in-built timeout for testpoints opened with DTT - if the measurement is inactive for some time (unsure how much exactly but something like 5mins), the testpoint is automatically closed.

  13241   Tue Aug 22 16:56:54 2017 johannesSummaryGeneralAS laser existing components inventory

I surveyed the lab today to see what we may need to buy for the AS laser setup.

We have:

NPRO 200 mW + Driver

Faraday Isolator from cabinet

ISOMET Model 1201E: This is a free space AOM I found in the modulator cabinet. It needs to be driven at 40MHz (to be confirmed) with ~6W of electrical power. For a 500 micron beam it can allegedly achieve rise times of '93' [units not specified, could this be nanoseconds?]. I did not find a dedicated driver for it, however there was a 5W minicircuits amplifier ZHL-5W-1 in the RF cabinet and a switch ZSDR-230, which has a typical switch time of 2 microseconds, however I'm not sure how this translates to rise/fall times of the deflected power. It seems we have everything to set this up, so we'll by the end of the week if we can use a combination of these things or if we need to buy additional driver electronics.

New Focus model 4004 broadband phase modulator which is labeled as dusty, and in fact quite dirty when looking through. We should attempt to clean this thing and maybe we can use it here or at the ends.

Probably all the optics we need for the PSL table setup.

 

We need:

Beat PD: How about one of these: EOT ET-3000A? I didn't find a broadband PD for the beat with the PSL

Fiber Stuff: coupler & polarization maintaining fiber 20m & collimator. There are a couple options here, which we can discuss in the meeting.

Faraday Isolator: If we want to inject P-polarization. If S is okay we can use a polarizing plate beamsplitter instead.

Possibly some large lenses for mode-matching to IFO (TBD)

 

 

  13240   Tue Aug 22 15:40:06 2017 gautamUpdateComputersOld frames accessible again

[jamie, gautam]

I had some trouble getting the daqd processes up and running again using Jamie's instructions.

With Jamie's help however, they are back up and running now. The problem was that the mx infrastructure didn't come back up on its own. So prior to running sudo systemctl restart daqd_*, Jamie ran sudo systemctl start mx. This seems to have done the trick.

c1iscey was still showing red fields on the CDS overview screen so Jamie did a soft reboot. The machine came back up cleanly, so I restarted all the models. But the indicator lights were still red. Apparently the mx processes weren't running on c1iscey. The way to fix this is to run sudo systemctl start mx_stream. Now everything is green.

Now we are going to work on trying the fix Rolf suggested on c1iscex.

Quote:

It turns out the problem was just a bent pin on the SCSI cable, likely from having to stretch things a bit to reach optimus from the RAID unit.frown

I hooked it up to megatron, and it was automatically recognized and mounted. yes

I had to turn off the new FB machine and remove it from the rack to be able to access megatron though, since it was just sitting on top. FB needs a rail to sit on!

At a cursory glance, the filesystem appears intact. I have copied over the achived DRFPMI frame files to my user directory for now, and Gautam is going to look into getting those permanently stored on the LDAS copy of 40m frames, so that we can have some redundancy.

Also, during this time, one of the HDDs in the RAID unit failed its SMART tests, so the RAID unit wanted it replaced. There were some spare drives in a little box directly under the unit, so I've installed one and am currently incorporating it back into the RAID.

There are two more backup drives in the box. We're running a RAID 5 configuration, so we can only lose one drive at a time before data is lost.

 

  13239   Tue Aug 22 15:17:19 2017 ericqUpdateComputersOld frames accessible again

It turns out the problem was just a bent pin on the SCSI cable, likely from having to stretch things a bit to reach optimus from the RAID unit.frown

I hooked it up to megatron, and it was automatically recognized and mounted. yes

I had to turn off the new FB machine and remove it from the rack to be able to access megatron though, since it was just sitting on top. FB needs a rail to sit on!

At a cursory glance, the filesystem appears intact. I have copied over the achived DRFPMI frame files to my user directory for now, and Gautam is going to look into getting those permanently stored on the LDAS copy of 40m frames, so that we can have some redundancy.

Also, during this time, one of the HDDs in the RAID unit failed its SMART tests, so the RAID unit wanted it replaced. There were some spare drives in a little box directly under the unit, so I've installed one and am currently incorporating it back into the RAID.

There are two more backup drives in the box. We're running a RAID 5 configuration, so we can only lose one drive at a time before data is lost.

ELOG V3.1.3-